OpenStudy (anonymous):

68. The slope of the line through the points (–3, 2) and (–3, 6) is: (not sure how to do these, can you help by showing work and i'll do the other ones on my own) thx

OpenStudy (anonymous):

\[s = (y _{2} - y _{1}) / (x _{2} - x _{1}) \] s = (6 - 2)/ (-3 -(-3)) s = (4)/(0) s = undefined because you can not divide by zero.